OpenStudy (michealsmith418625):

y=700−22(x) GIVING MEDALS

OpenStudy (dinamix):

so what your question

OpenStudy (michealsmith418625):

how many beams will remain if she builds six floors

OpenStudy (michealsmith418625):

x=6

OpenStudy (michealsmith418625):

the answer is 568

OpenStudy (dinamix):

f(6) = 700-(22)(6) = 700-132 =568 just put x and remplace it in equation @michealsmith418625

OpenStudy (michealsmith418625):

how do i give medals

OpenStudy (dinamix):

best response click lol

OpenStudy (michealsmith418625):

okay thanks lol

OpenStudy (jjcoolja12):

Let's solve for x. y=700−22x Step 1: Flip the equation. −22x+700=y Step 2: Add -700 to both sides. −22x+700+−700=y+−700 −22x=y−700 Step 3: Divide both sides by -22. −22x −22 = y−700 −22 x= −1 22 y+ 350 11 Answer: x= −1 22 y+ 350 11